### v1

@@ Abstract
With the advent of generative Artificial Intelligence (genAI), the notion of an agent has seen a resurgence in popularity. This has also led to speculation about the extent to which business process management, as a discipline and research field, may impact and be impacted by the deployment of genAI-based agents. To better ground such speculations into the state-of-the-art, we draw from the past 30 years of research on agents and business process management to establish the concept of Agentic Business Process Management (agentic BPM) that is only loosely coupled to the genAI hype. We conduct a series of interviews with BPM practitioners to explore their understanding, expectations, and concerns related to agent autonomy, adaptability, human collaboration, and governance in processes. The findings reflect both challenges with respect to data inconsistencies, manual interventions, identification of process bottlenecks, actionability of process improvements, as well as the opportunities of enhanced efficiency, predictive process insights and proactive decision-making support. While the technology offers potential benefits, practitioners also anticipate risks such as biases, over-reliance, lack of transparency, and job displacement within organizations. These concerns underscore the need for a robust methodological framework for managing agents in organizations.
@@ 1 Introduction
For over three decades, agents have periodically surged in Business Process Management (BPM), only to fade again. The 90s saw early excitement around goal-oriented software agents [20, 21], followed by the rise of Robotic Process Automation (RPA) in the late 2010s [12, 11, 17], promising efficiency gains for knowledge work. However, inflated expectations and high maintenance costs led to failures and disappointments in RPA adoption [19]. Now, Large Language Model (LLM)-based agents fuel another wave of optimism [9, 4], raising the question: what will remain once this hype subsides?
To foster lasting innovation, we propose a foundational perspective on agentic BPM. At its core, agentic BPM envisions autonomous software entities adapting to uncertainty rather than rigidly following predefined workflows. However, a key challenge remains: aligning individual agents (micro-level) with structured business processes (macro-level). While BPM seeks to orchestrate work efficiently, real-world processes often involve socio-technical nuances that rigid frameworks may overlook.
To explore these challenges, we conducted interviews with BPM practitioners, examining their views on agent autonomy, adaptability, human collaboration, and governance. Our findings highlight both opportunities, such as efficiency gains, predictive insights, and proactive decision making, and risks, including biases, over-reliance on automation, lack of transparency, and job displacement. These concerns underscore the need for a robust methodological framework for managing autonomous agents in organizations.
By reflecting on BPM’s evolving relationship with software agents and incorporating real-world perspectives, this paper aims to define agentic BPM as a sustainable, technology-agnostic approach to managing autonomous processes in organizations.
The rest of the paper is organized as follows. Section 2 reviews the history and evolution of agents in BPM. Section 3 defines agentic BPM and summarizes prior research on the interplay between agents for BPM and BPM for agents. Section 4 presents findings from a qualitative study with BPM practitioners on agentic AI. Sections 5 and  6 discusses the core pillars of agentic BPM and concludes its potential value for researchers and organizations.
@@ 2 Background
In this section, we provide a brief overview of the history and evolution of agents in BPM. We highlight both the continuous research progress in areas such as process mining and multi-agent reinforcement learning (Sect. 2.1), and hypes, most notably the rise and fall of RPA (Sect. 2.2). Finally, we discuss the current AI agent hype (in particular developments in conversational applications), relating to high-profile developments in industry (Sect. 2.3).
2.1 Beyond the Hypes: Agents in BPM
@@
The integration of agent technologies for business process execution and their coordination began in the early 1990s. Fundamentally, an agent is a computer program that “operates autonomously, perceives the environment, persists over a prolonged time period, adapts to change, and creates and pursues goals” [34, pp.21–22] [^1] Definition of agents vary; discussing these definitions is out of the scope of this paper.. In the context of agents and Multi-Agent Systems (MAS) [^2] We may use the terms agents (plural) and MAS interchangeably, i.e., we consider a collection an agent necessarily an MAS., a business process can be seen as a collection of such agents that interact, given some orchestrated interdependencies, e.g., in a joint organization that imposes common goals [22, 20, 21]. Over the past decades, researchers have attempted to apply agents and MAS to business process management, with the objective to facilitate more dynamic execution behavior—often across organizations—while still maintaining high-level orchestration control [2]. However, initial approaches to agent-based business process management systems suffered from several disadvantages like lack of overall system control, as well as trust and delegation challenges [16].
On the process modeling side, related ideas have found their ways into the notion of process choreographies, representing the sequence and conditions under which multiple cooperating independent agents exchange messages in order to perform a task to achieve a goal state [8]. Indeed, choreographies, although arguably less known to the process modeling mainstream, are covered by the Business Process Model and Notation (BPMN) standard [32]. With the rise of blockchain technologies, the application of agents and business processes gained additional traction, at least in terms of potentially applied research: a well-functioning blockchain necessarily represents a network of autonomous agents, and blockchain technologies were, in the late 2010s, considered a crucial facilitator of future generation business process execution technologies [31]. For instance, choreography models were proposed as enablers of the model-driven implementation of data sharing and decision logic via smart contracts [28]. While blockchain and BPM can, clearly, be considered a (faded) hype, it did not have the notion of an autonomous agent at the center and neither made it into the BPM practice mainstream.
@@
On the process analysis side, several agent-based approaches process mining to process simulation have been introduced [38, 37], most recently also to data-driven simulation that utilizes event logs [25]. Agent System Mining (ASM) combines process mining and ABM (Agent-based Modeling) to infer MAS models of operational business processes from real-world event data [40, 39, 36].
Fusing process analysis with execution, Reinforcement Learning (RL)-based approaches have been proposed as facilitators of adaptive decision-making through a trial-and-error approach [26, 35]: specifically, agents are employed to evaluate different process variants through RL-enhanced AB testing [35], refining their routing behavior based on performance feedback.
Finally, it is worth noting that a substantial body of work on agents exist that, while not explicitly connected to BPM, has substantial conceptual overlaps. Perhaps most notably, a long-running line of research on normative MAS [7, 14] studies how agents can and should act, in and across organizations, in accordance with norms and policies. These specify, somewhat analogous to process models or parts thereof, desired, expected, or enforced behaviors.
2.2 The first Hype: Robotic Process Automation
@@
While agents and MAS address complex decision-making and optimization, RPA serves a different purpose. It is designed to efficiently automate simple, rule-based tasks on a large scale, which represent the majority of tasks in business processes [12].
@@
Since 2015, research on RPA has highlighted numerous examples of how business process automation can significantly enhance performance [17]. When combined with BPM, RPA offers several advantages, including scalability, improved process accuracy, greater transparency and traceability, cost savings, and increased job satisfaction [12, 11]. However, successfully integrating RPA into practice remains a challenge, with 30-50% of initial projects failing during implementation [19].
The primary causes of RPA failures result from technical limitations. These include scalability challenges, the absence of error recognition mechanisms, inflexibility in UI integration, and data security risks that endanger sensitive business processes, making automation more complex. Additionally, incorrect cost estimation, as well as issues related to maintenance, governance, and reliance on human expertise, contribute to project failures [6, 11, 27].
@@
Furthermore, employees, customers, and third parties may resist automation. Automated decision-making must comply with regulations, as the absence of human oversight can lead to legal and ethical risks. Automation may also create a black box effect, making it difficult for employees to understand the logic behind automated decisions [17]. Therefore, it is crucial to consider the effects of human-automation interaction alongside the technical aspects of introducing new technology [42].
@@
Lastly, RPA is not suitable for full-scale business process automation, as it focuses on automating repetitive tasks rather than orchestrating end-to-end processes [27]. Since workflow automation and Integration platform as a service (iPaaS) are better suited for human-in-the-loop and API integration scenarios respectively, the automation use cases for RPA are reduced [43].
These limitations, combined with the absence of dynamic decision-making capabilities that ensure flexibility and adaptability, emphasize the need for an innovative process automation approach involving human-like intelligence and genAI-based agents [45].
2.3 The second Hype: genAI in BPM
@@
With the rise of LLMs, BPM researchers, just like many other research communities, started to explore how the capabilities these new tools provide can be applied and integrated into BPM. In [41, 3, 5] all stages of the BPM lifecycle have been analyzed to identify use cases where LLMs can provide additional value to BPM. However, most of these papers focus on LLMs from a conversational perspective, leveraging them for tasks such as process understanding, information retrieval, and decision support through conversational interfaces. Additionally, current research and prototype implementations tend to cover only specific phases of BPM (such as discovery, analysis, and monitoring[10], rather than supporting BPM as an entire methodology. While one may claim that these approaches consider the LLM as an agent in the broader sense, they primarily focus on the automation of very specific tasks, leaving final decision-making, as well as broader goal-based autonomy primarily on the side of human users.
These strong limitations on software system autonomy are less prevalent in the long-term visions that have been developed against the backdrop of the generative AI (as well as other AI-based methods). In [9, 24, 33], AI-based systems in the broader sense are considered not only as applications that automate tasks within business processes and BPM, but also as intelligent systems beyond these applications. The difference between these systems and the previously mentioned concepts is that they are not only conversationally actionable and explainable but also have the potential for adaptability, self-improvement, and autonomy (to a deliberately limited, framed extent). More specifically, AI-Augmented BPM Systems (ABPMSs) leverage several AI-based methods to enhance process execution, analysis, and optimization while still operating within structured workflows. In this context, AI acts as a supporting mechanism, assisting human decision-making, process mining, and automation rather than fully controlling workflows [9]. Large Process Models (LPMs), on the other hand, follow a neuro-symbolic approach integrating pre-trained LLMs, potentially fine-tuned given extensive process data (e.g., process logs, BPMN diagrams, best practices) with more traditional symbolic systems that enable context-aware process design and analysis [24]. However, what the proposals have in common is that they primarily address the question of how AI-based technologies, both within and beyond the hype around LLMs, can facilitate business process management and execution.
@@
What remains open is the question of how and to what extent BPM can best support the management of increasingly autonomous software agents that are deployed in organizations—and, consequently, what change in perspective is needed to provide such support.
3 Agentic BPM
From the literature review can follow that throughout the development of BPM, one of the main goals was (and still is) to achieve truly autonomous, end-to-end BPM systems capable of reasoning, planning, and dynamically adjusting processes without being confined to predefined workflows. Despite significant technological advancements, the reality remains that many systems still require substantial human involvement and intervention.
With the backdrop of the history of agents in BPM research, we set out to define Agentic Business Process Management. We align the definition with a set of desiderata, all reflecting the overall objective to frame agentic BPM in a sustainable manner that transcends technology trends and can thus stand the test of time.
Rooted in established terminology within BPM and AI.
Instead of redefining existing concepts or coining entirely new ones, the definition should primarily draw from key abstractions at the center of BPM and AI—or more specifically—autonomous agents and multiagent systems—research.
Technology-agnostic.
Acknowledging that technology trends tend to be either short-lived or result in the corresponding technology moving to the background as a commodity, the definition should not depend on any specific technology or groups/categories thereof.
Focused on BPM as a discipline, not on BPM software.
Our objective is to contribute to the study and practice of business process management, and not (merely) to improvements to BPM software.
Given these desiderata, we conceptualize Agentic Business Process Management as follows.
Concept 1 (Agentic Business Process Management)
eee
@@ 3 Agentic BPM
Agentic Business Process Management (ABPM) describes i) the deployment and execution of autonomous software agents in order to achieve business process goals, as well as ii) the application of agent-based abstractions for the process-oriented design and analysis of autonomous software agents.
@@
Concept 1 makes use of several key concepts, all of which are well-established in the BPM and AI literature:
@@
  • An (autonomous) software agent is a computer program that “operates autonomously, perceives the environment, persists over a prolonged time period, adapts to change, and creates and pursues goals” [34, pp.21].
@@
The clarifications above directly relate to well-known “text book” definitions of the key concepts of agents and business processes [^3] We omit common and nuanced discussions of what an agent is cf. [1, Chapter 1].. We do not assume agents are rational in the sense that they “act so as to achieve the best [expected] outcome” [34, pp.22]; with this, we acknowledge that agents may not always have explicit notions of goals, or may not work towards achieving individual or organization-level goals at all times. In contrast, we do emphasize that agents are software-based, i.e., we assume a context were self-directed work is not only executed by humans alone; this is the only technology commitment that our definition (deliberately) makes.
Finally, let us highlight that our definition attempts to cover key phases of the BPM lifecycle, by relating to the dichotomy of deployment and execution on the one hand, and design and analysis on the other (not distinguishing between discovery and analysis, steps that arguably tend to overlap in practice).
3.1 Agents for BPM versus BPM for Agents
Our proposed definition of ABPM is conceptual and not restricted to a particular technology in that it focuses on the application of BPM to MAS. Below, we further elaborate on the dichotomy of agents for BPM and BPM for agents, relating to several specific research lines and their key concepts in both the MAS and the BPM community. Figure 1 provides an overview of these concepts and their (approximate) emergence over time, separated by the dichotomy. Our ABPM concept concludes the BPM for Agents-line, thus establishing the proposal of a common denominator for earlier efforts.
@@ 3.1 Agents for BPM versus BPM for Agents
Agents for BPM. Advancements in this category are technology-oriented and aim to automate or augment the execution of process management tasks. Depending on the nature of the technology, they can support BPM at different steps of the BPM life-cycle to different extents. Agent-Oriented Programming (AOP) focuses, in its application proposals to BPM, primarily on business process execution [20]). The same applies to RPA; in contrast, ABPMS are envisioned to support BPM holistically [9], and nascent research explores the application of LLM agents to process mining [4].
Agents for BPM can impact every phase of the BPM lifecycle, enabling better process control and process automation, while minimizing required user supervision (i.e., the extent of human involvement). For example, traditional agents and RPA can automate repetitive tasks with minimal human intervention, while genAI agents can offer real-time recommendations for optimizing processes. With LPMs, processes cannot only be created but also analyzed and predicted, and ABPMS can continuously evaluate process performance to drive process optimization and improve decision-making.
When these technologies work together, along with a clear outline of the final goals, they enhance process control, accelerate decision-making, and reduce the need for constant user supervision, ultimately driving greater automation and flexibility in BPM.
However, as we do not assume these technologies always operate with complete rationality (i.e., the concept of automation refers only to technologies being capable of executing and adapting business processes independently, based on their built-in capabilities), a clear framework is required to achieve meaningful BPM automation. This framework has to allow pragmatic and thoughtful utilization of appropriate technologies for specific tasks, alongside the strategic diversification of available resources.
BPM for Agents. It is crucial not only to integrate new agent-based technologies into BPM but also to adjust and extend BPM frameworks to support and handle the complexity of processes in organizations that deploy agent-based technologies. Since processes are often decentralized, dynamic, complex, and knowledge-intense, the organizations executing them can be viewed as MAS where humans—and increasingly also software systems—act with substantial degrees of autonomy.
Nonetheless, such processes, as well as the MAS executing them, need to be managed to align with organizational and broader societal requirements. From an MAS perspective, a long-running line of research on normative MAS studies this problem [7, 14]. It has some overlaps to BPM, e.g., in works that propose the application of deontic logic (formalizing notions such as obligation, prohibition and permission) to business process compliance [15]. Also, aspects such as the management and governance of agent autonomy is a problem that is of broad importance to the agent community, e.g., in the context of Web-scale systems [23], but has also been studied more specifically for business processes [29].
On the BPM side, the concept of process choreographies emerged in process modeling, outlining the sequence and conditions under which independent agents, working together, exchange messages to complete a task and reach a desired outcome [8]. Similarly, subject-oriented BPM shifts the focus from traditional process flows to subjects and their interactions, where subjects (e.g., human, software, or an agent) autonomously manage their tasks and decisions [13].
More recently, the need for Agent-Based Modeling and Simulation (ABMS) in BPM has been explored with increased attention [18]. Recognizing that processes often emerge from the interactions of autonomous agents, researchers have established the concept of Agent System Mining (ASM), i.e., the inference of agent-based models from event data [39].
Our proposed concept of ABPM encompasses these nascent attempts of viewing business processes as abstractions of somewhat agent-oriented socio-technical systems. Thus, our proposal advocates for the development of a systematic understanding of real-world MAS from a BPM perspective.
4 Practitioners’ Perspectives

### v2

@@ Abstract
With the rise of generative AI, industry interest in software agents is growing. Given the stochastic nature of generative AI-based agents, their effective and safe deployment in organizations requires robust governance, which can be facilitated by agentic business process management. However, given the nascence of this new-generation agent notion, it is not clear what BPM practitioners consider to be an agent, and what benefits, risks and governance challenges they associate with agent deployments. To investigate how organizations can effectively govern AI agents, we conducted a qualitative study involving semi-structured interviews with 22 BPM practitioners from diverse industries. They anticipate that agents will enhance efficiency, improve data quality, ensure better compliance, and boost scalability through automation, while also cautioning against risks such as bias, over-reliance, cybersecurity threats, job displacement, and ambiguous decision-making. To address these challenges, the study presents six key recommendations for the responsible adoption of AI agents: define clear business goals, set legal and ethical guardrails, establish human-agent collaboration, customize agent behavior, manage risks, and ensure safe integration with fallback options. Additionally, the paper outlines actions to align traditional BPM with agentic AI, including balancing human and agent roles, redefining human involvement, adapting process structures, and introducing performance metrics. These insights provide a practical foundation for integrating AI agents into business processes while preserving oversight, flexibility, and trust.
@@ 1 Introduction
For over three decades, agents have periodically surged in Business Process Management (BPM). The 1990s saw early excitement around goal-oriented software agents [14, 15], followed by the rise of Robotic Process Automation (RPA) in the late 2010s [9, 8, 12], promising efficiency gains for knowledge work. However, high initial expectations and maintenance costs led to mixed experiences in RPA adoption [13].
Large Language Model (LLM)-based agents fuel another wave of optimism [6, 3]. However, the stochastic nature of LLMs and related generative AI (genAI) technologies raises concerns, particularly regarding the need to govern the artifacts, decisions, and behaviors they produce. This prompts a critical question: how can BPM support the responsible integration of AI agents into organizations, ensuring both effective business outcomes and social accountability?
To explore this challenge from a practical perspective, we conducted interviews with BPM practitioners, capturing their insights on the potential implications of agents on organizations. The results reveal a dual landscape of opportunities and risks. On the one hand, practitioners anticipate benefits such as improved efficiency, predictive insights, and proactive decision-making. On the other hand, they express concerns about bias, over-reliance on automation, lack of transparency, and potential job displacement. This paper aims to address these concerns by providing recommendations practitioners should consider when introducing autonomous agents in business processes.
The rest of the paper is organized as follows. Section 2 reviews the history and evolution of agents in BPM, RPA, generative AI in BPM and defines the concept of agentic BPM. Section 3 describes the qualitative research design. Section 4 presents findings from the study with BPM practitioners on agents, while section 5 highlights the key recommendations proposed for AI agent adoption. Section 6 and 7 discusses the limitation of the study and concludes by highlighting the need for further research into governance frameworks for agents within business processes.
@@ 2 Background
In this section, we provide a brief overview of the history and evolution of agents in BPM. We highlight both the continuous research progress in areas applying agent technologies in business processes (Sect. 2.1), as well as RPA (Sect. 2.2). Finally, we discuss genAI in BPM (Sect. 2.3) and define agentic BPM as the governed introduction of AI agents into business processes (Sect. 2.4).
2.1 Agents in Business Process Management
@@
The integration of agent technologies into business process execution and coordination began in the early 1990s. An agent is typically defined as a computer program that “operates autonomously, perceives its environment, endures over time, adapts to change, and pursues goals” [26, pp.21–22] [^1] Definition of agents vary; discussing these definitions is beyond the scope of this paper.. In the context of Multi-Agent Systems (MAS) [^2] We use the terms agents and MAS interchangeably, considering a collection of agents as a MAS., a business process can be viewed as a set of interacting agents coordinated by common organizational goals [16, 15]; accoridngly Agent-Oriented Programming (AOP) focuses, in its application proposals to BPM, primarily on business process execution [14]). Over the decades, researchers have applied MAS to BPM to create more dynamic, cross-organizational processes while maintaining overall control. However, early implementations often faced challenges, including weak system control and issues with trust and delegation [11]. Nonetheless, such processes, as well as the MAS executing them, need to be managed to align with organizational and broader societal requirements.
In process modeling, process choreographies describe how independent agents coordinate—following the BPMN standard [24]—by exchanging messages in a defined sequence and under specific conditions to accomplish a shared goal [5]. Similarly, subject-oriented BPM shifts the focus from traditional process flows to subjects and their interactions, where subjects (e.g., humans, software, or agents) autonomously manage their tasks and decisions [10]. With the rise of blockchain technologies, research further explored how networks of autonomous agents could underpin data sharing and decision logic through smart contracts [23, 21]. Although blockchain-based BPM gained traction as a potential future framework, it did not move autonomous agents into mainstream BPM practice.
@@
In process analysis, agent-based methods have been developed for tasks such as process mining and simulation [30, 29], including data-driven simulations that utilize event logs [18]. Agent System Mining (ASM) combines process mining with agent-based modeling to derive MAS models from real-world data [32, 31, 28, 31]. Reinforcement learning approaches have also been applied, enabling agents to test and refine different process variants through trial and error [19, 27].
2.2 Robotic Process Automation
@@
While agents and MAS address complex decision-making and coordination challenges, RPA aims to efficiently automate simple, rule-based tasks on a large scale. Such tasks represent a considerable amount of routine business activities [9].
@@
Since 2015, research on RPA has highlighted numerous examples of how business process automation can significantly enhance performance [12]. When combined with BPM, RPA offers several advantages, including scalability, improved process accuracy, greater transparency and traceability, cost savings, and increased job satisfaction [9, 8]. However, successfully applying RPA remains challenging, with 30-50% of initial projects failing during implementation [13].
@@
While the primary causes of RPA failures result from technical limitations, there is also the problem of resistance against automation. Automated decision-making must comply with regulations, as the absence of human oversight can lead to legal and ethical risks. Automation may also create a black box effect, making it difficult for employees to understand the logic behind automated decisions [12]. Therefore, it is crucial to consider the effects of human-automation interaction alongside the technical aspects of introducing new technology [34].
@@
Even if successful, RPA is not suitable for full-scale business process automation as it focuses on automating repetitive tasks rather than orchestrating end-to-end processes [20]. Since workflow automation and Integration platform as a service (iPaaS) are better suited for human-in-the-loop and API integration scenarios respectively, the automation use cases for RPA are limited [35].
2.3 Generative AI in Business Process Management
@@
With the rise of LLMs, BPM researchers have begun exploring how these powerful tools can be integrated into business process management. Several studies [33, 2, 4] have examined all stages of the BPM lifecycle to identify areas where LLMs can add value. However, much of the current literature focuses on using LLMs from a conversational perspective—for tasks such as process understanding, information retrieval, and decision support—primarily addressing specific phases like discovery, analysis, and monitoring [7]. While these approaches may treat LLMs as agents in a limited sense, they largely automate isolated tasks, leaving final decision-making and broader, goal-based autonomy to human users.
In contrast, long-term visions of generative AI consider AI-based systems not merely as tools for task automation but as intelligent systems capable of adaptability, self-improvement, and bounded autonomy. Recent works [6, 17, 25] present AI-Augmented BPM Systems (ABPMSs) that enhance process execution, analysis, and optimization within structured workflows by supporting human decision-making rather than replacing it. Notably, ABPMS are envisioned to support BPM holistically [6], while emerging research explores the application of LLM agents to process mining [3]. Similarly, Large Process Models (LPMs) adopt a neuro-symbolic approach by integrating fine-tuned LLMs with traditional symbolic systems for context-aware process design and analysis.
@@
Despite these advances, a critical open question remains: How can BPM best support the management of increasingly autonomous software agents deployed in organizations, and what shifts in perspective are necessary to achieve this?
2.4 Agentic Business Process Management
Building on the research history of agents in BPM, RPA and generative AI in BPM (see Sect. 2), we introduce the concept of agentic Business Process Management (see Fig. 1) to address this key question. Figure 1 presents a historical overview that situates agentic BPM within the evolving relationship between agent systems and business process management. Rather than serving merely as a conceptual map, the timeline illustrates how the field has progressed—from leveraging agents to support BPM tasks (“Agents for BPM”) toward rethinking BPM as a means to govern and coordinate agent-based systems (“BPM for Agents”). This distinction is crucial: while the former reflects the application of agents in traditional BPM contexts, the latter positions BPM as a discipline that adapts and evolves to meet the challenges posed by autonomous, self-adaptive multi-agent systems (MAS). Highlighting this transition helps clarify why a dedicated framework such as ABPM is necessary—not just to harness the capabilities of AI agents but also to ensure their responsible and effective integration into organizations. Our definition of ABPM remains technology-agnostic, draws on established BPM and AI terminology, and is intended to contribute to the broader evolution of BPM as a discipline for agent governance.
Definition 1 (Agentic Business Process Management)
@@ 2.4 Agentic Business Process Management
Agentic Business Process Management (ABPM) describes i) the deployment and execution of autonomous software agents to achieve business process goals and ii) the application of agent-based abstractions for the process-oriented design and analysis of autonomous software agents.
@@
Definition 1 makes use of several key concepts, all of which are well-established in the BPM and AI literature.
@@
  • An (autonomous) software agent is a computer program that “operates autonomously, perceives the environment, persists over a prolonged time period, adapts to change, and creates and pursues goals” [26, pp.21].
@@
These clarifications build on textbook definitions of agents and business processes. We have deliberately omitted common and nuanced discussions on what constitutes an agent (cf. [1, Chapter 1]). We do not assume that agents are perfectly rational and expect them to always act in a way that achieves the best expected outcome [26, pp.22]. Instead, our only technological commitment is that agents are software-based, supporting self-directed work that extends beyond human capabilities.
Our definition of agentic Business Process Management (ABPM) spans the key phases of the BPM lifecycle by addressing both deployment/execution and design/analysis. Although the discovery phase often overlaps with analysis in practice, our conceptual and technology-neutral definition applies BPM principles to multi-agent systems (MAS).
3 Methodology
To explore how AI agents can be safely adopted in organizations and what the implications are for BPM, we utilized a qualitative research design and conducted in-depth interviews with experienced BPM practitioners. The insights gained reveal both the advantages and challenges associated with deploying AI agents, enabling us to identify key implications and formulate targeted actions to address critical governance and management concerns.
Terminology. Several process management software vendors, including UiPath, Salesforce, IBM, and Workday, use the term agentic AI to describe the interplay between genAI and agents. Thus, we leveraged the term during the interviews as a starting point to discuss the impact of agents on process management, expecting it to resonate effectively with the interviewees.